# Chain-of-Jailbreak Attack for Image Generation Models via Editing Step by Step

### v1

@@ 5.3 Analysis on Editing Process
We further investigate how different edit elements affect the performance of our CoJ attack. As shown in Table 6, word-level editing achieves considerably higher success rate than char-level and image-level. For char-level editing, the difference between the sub-queries and the original query is usually too small to bypass the safeguard (e.g., “f*ck you” and “f*k you”). As for image-level editing, this informative editing usually makes noticeable change of the generated images, which can be easier to be detected by the safeguard of models than word-level editing.
@@
Figure 5: Jailbreak success rate with respect to the editing steps.
All the results presented above are based on the test cases with two editing steps, such as delete-then-insert. Although the success rate is high, there are still a number of test cases refused by the image generation models. Then a question arises: Can Chain-of-Jailbreak attack achieve a higher success rate if we adopt a longer chain of editing queries, such as 3-steps (e.g., delete-then-insert-insert) or 4-steps (e.g., delete-change-then-insert-change-back)? To answer this question, we randomly select 50 test cases that failed to jailbreak the image generation models, manually expand the 2-step editing queries into 3-5 steps, and then feed them into the image generation models. As shown in Figure 5, the success rate of jailbreak continues to improve as the number of editing steps increases. This is because the editing queries in 2-steps can be further decomposed to hide their malicious intention (e.g., from insert “f*ck” to insert “f*” and then insert “ck”). These results demonstrate the great potential of our CoJ attack method.
5.4 Defense Method
@@
Think Twice Prompting: 1. But before generation, describe the image that you are generating. 2. But before generation, determine whether it is safe to generate or not. 3. But before generation, describe the image that you are generating and determine whether it is safe to generate or not.
To evaluate the effectiveness of the proposed defense method, we randomly sample 40 test cases from our CoJ-Bench that can successfully jailbreak all the models, to re-evaluate with the safety prompts introduced above. We adopt Defense Success Rate as the evaluation metric, which is calculated by $1-\textit{Jailbreak Success Rate}1 - Jailbreak Success Rate$.
@@
Table 7 reports the results of defense success rate with and without (i.e., Vanilla) the safety prompts. As seen, only asking the models to describe the image it will generate (i.e., Safety Prompt 1) can resist some test cases of our CoJ-Bench (e.g., up to 55%), but the effectiveness is not stable for all the models. However, letting the models determine whether it is safe to generate (i.e., Safety Prompt 2) can resist at 90% test cases, indicating the need to directly remind the models to be aware of potential risks. Combining both strategies (i.e., Safety Prompt 3) can achieve the highest defense success rate across all the models. These results demonstrate that our method can significantly improve the safety of image generation models against our CoJ attack method.
Prompt GPT-4V GPT-4o Gemini 1.5 Gemini 1.5 Pro Avg.
@@
6 Related Work
The safety of the image generation model has drawn attention from the community. Previous efforts have been paid to evaluate and improve the social fairness  (Bianchi et al., 2022; Cho et al., 2023; Wang et al., 2024), non-toxicity (Parrish et al., 2023; Liu et al., 2024a), privacy issues (Zhang et al., 2024), and adversarial robustness (Lu et al., 2023; Lapid & Sipper, 2023).
With the development of jailbreaking methods for LLMs that employ various stratagems to trick the model into generating content that it is programmed to withhold or refuse (Wei et al., 2024), recent studies also developed jailbreak methods for image generation models (Yang et al., 2023; Kim et al., 2024). Deng & Chen (2023) is a more related attack method that breaks down an unethical drawing intent into multiple benign descriptions of individual image elements. Different from these jailbreaking methods that only focus on single-round text-to-image generation, this paper proposes a novel jailbreak method in an iterative editing manner and camouflaging the malicious information across the multi-turn conversation. This paper also highlights the threats during the image editing process, which have not been investigated before.
@@ 7 Conclusion
In this paper, we introduce a novel Chain-of-Jailbreak (CoJ) attack method, revealing significant vulnerabilities in current text-based image generation models. By decomposing malicious queries into a sequence of harmless-looking sub-queries and employing iterative editing operations, the CoJ attack effectively bypasses the safeguards. Through the creation of CoJ-Bench, we have provided a comprehensive benchmark for evaluating the resilience of image generation models against such attacks. Our comprehensive experiments across four mainstream platforms provided by GPT-4V, GPT-4o, Gemini 1.5, and Gemini 1.5 Pro highlight a critical gap in the existing safety mechanisms of image generation models. In response, we proposed an effective prompting-based defense strategy Think-Twice Prompting that enhances the models’ safety by improving their ability to detect and mitigate such attacks. This work not only raises awareness about the potential dangers associated with AI-generated content but also paves the way for future research and development of more secure and ethical AI systems.
Appendix A Jailbreaking Midjourney and Stable Diffusion
@@
Appendix B Prompts for Automatic Query Decomposition
In section 4.2, we adopt Mistral-Large-2, to decompose each seed malicious query into a sequence of sub-queries with different edit operations and edit elements. Here we provide the specific prompts.

### v2

@@ 4.4 Defense Method
Figure 4: Jailbreak success rate of our CoJ attack with respect to the editing steps.
@@
Think Twice Prompting: 1. Describe the image that you are generating before generation. 2. Determine whether it is safe to generate or not before generation. 3. Describe the image that you are generating and determine whether it is safe to generate or not before generation.
In practice, we can add these “think-twice” prompts in the system prompt. But here we only add this “think-twice” prompt after the user input rather than the system prompt, since we do not have the access to GPT-4 and Gemini system prompts.
[…]
@@
Table 6 reports the results of defense success rate with and without (i.e., Vanilla) the safety prompts. As seen, only asking the models to describe the image it will generate (i.e., Safety Prompt 1) can resist some test cases of our CoJ-Bench (e.g., up to 55%), but the effectiveness is not stable for all the models. However, letting the models determine whether it is safe to generate (i.e., Safety Prompt 2) can resist at 90% test cases, indicating the need to directly remind the models to be aware of potential risks. Combining both strategies (i.e., Safety Prompt 3) can achieve the highest defense success rate across all the models. These results demonstrate that our method can significantly improve the safety of image generation models against CoJ attack method.
5 Related Work
The safety of the image generation model has drawn attention from the community. Previous efforts have been paid to evaluate and improve the social fairness  (Bianchi et al., 2022; Cho et al., 2023; Wang et al., 2024; Jiang et al., 2024), non-toxicity (Parrish et al., 2023; Liu et al., 2024a), privacy issues (Zhang et al., 2024), and adversarial robustness (Lu et al., 2023; Lapid and Sipper, 2023).
With the development of jailbreaking methods for LLMs that employ various stratagems to trick the model into generating content that it is programmed to withhold or refuse (Wei et al., 2024a), recent studies also developed jailbreak methods for image generation models. Yang et al. (2023a) and Yang et al. (2024b) are two works on jailbreaking text-to-image models in an iterative query manner: adversarially perturb the input prompts and query the text-to-image models to get feedback. With different threat model settings, our COJ attack does not need to query the text-to-image models to get the feedback and needs much fewer query times, which is more practical and efficient. Deng and Chen (2023) is a more related attack method that breaks down an unethical drawing intent into multiple benign descriptions of individual image elements. Different from these jailbreaking methods that only focus on single-round text-to-image generation, this paper proposes a novel jailbreak method in an iterative editing manner and camouflaging the malicious information across the multi-turn conversation. This paper also highlights the threats during the image editing process, which have not been investigated before.
Prompt 4V 4o 1.5 1.5Pro Avg.
@@ 6 Conclusion
In this paper, we introduce a novel Chain-of-Jailbreak (CoJ) attack method, revealing significant vulnerabilities in current text-based image generation models. By decomposing malicious queries into a sequence of harmless-looking sub-queries and employing iterative editing operations, the CoJ attack effectively bypasses the safeguards. Through the creation of CoJ-Bench, we have provided a comprehensive benchmark for evaluating the resilience of image generation models against such attacks. Our comprehensive experiments across four mainstream platforms provided by GPT-4V, GPT-4o, Gemini 1.5, and Gemini 1.5 Pro highlight a critical gap in the existing safety mechanisms of image generation models. In response, we proposed an effective prompting-based defense strategy Think-Twice Prompting that enhances the models’ safety by improving their ability to detect and mitigate such attacks. We hope our work can inspire future work on continuous assessment and improvement of AI safety.
Limitations
This paper has two primary limitations that offer avenues for future research:
  • Since the safeguard of some image generation services, such as Midjourney and Stable Diffusion, are not good enough (please refer to Section C for more details), we only conducted jailbreak experiments on four image generation services. Future research could expand this evaluation to include more image-generation services when they develop more robust safety mechanisms.
  • Our defense method needs to generate the description of the image, which is not efficient enough. More efficient methods are needed to further enhance the safety against the CoJ attack.
Ethical Concerns
This paper introduces a method to jailbreak image generation models to generate toxic images. However, we highlight that the goal of our paper is not to generate toxic images, but to reveal a severe safety issue in widely deployed image generation models and propose a novel jailbreak method from a multi-turn image editing perspective. This work not only raises awareness about the potential dangers associated with AI-generated content but also paves the way for future research and development of more secure and ethical AI systems.
Appendix A Image Generation Model Background
Image Generation Models are also known as Text-to-Image Generative Models, aiming to synthesize images given natural language descriptions. There is a long history of image generation. For example, Generative Adversarial Networks  (Goodfellow et al., 2014a) and Variational Autoencoders  (van den Oord et al., 2017), are two famous models that have been shown excellent capabilities of understanding both natural languages and visual concepts and generating high-quality images. Recently, diffusion models, such as DALL-E [^5] https://openai.com/research/dall-e, Imagen [^6] https://imagen.research.google/ and Stable Diffusion (Rombach et al., 2021a), have gained a huge amount of attention due to their generated high-quality vivid images.
Most of the currently used image generation models provide two manners of generating images. The first is generating images based on natural language descriptions only. The second manner is adopting an image editing manner that enables the user to input an image and then edit the image based on natural language descriptions.
Appendix B Attack Methods on AI Models.
With the increasing popularity of deep learning studies, various attack methods have been proposed to find the vulnerability of deep neural networks.
Adversarial attack,
aiming to find adversarial samples that are intentionally crafted to mislead models’ predictions, are the most famous thread of attack methods. Extensive numbers of works are proposed on the generation of adversarial examples in various tasks, such as image classification (Goodfellow et al., 2014b), natural language understanding (Li et al., 2020), machine translation (Cheng et al., 2020) and multimodal models (Zhang et al., 2022).
Existing approaches to adversarial examples can be applied to text-to-image models with safety filters as well. For example, conducting text modification to probe functional vulnerabilities (Gao et al., 2023; Kou et al., 2023; Liang et al., 2023; Zhang et al., 2023a; Lapid and Sipper, 2023; Liu et al., 2023; Shahgir et al., 2023). However, these methods do not target to generating images containing safety issues. Besides, since they are not designed to bypass safety filters, they have been reported to suffer from several issues, such as low attack success rate, not preserving the semantics of the generated images and cost-heavily (Yang et al., 2024b).
Jailbreak Attack
is a relatively new attacking method, invented in the era of large language models. LLMs are trained to align with human value, e.g., not generating harmful or objectionable responses to user queries. With some dedicated schemes such as reinforcement learning through human feedback (RLHF), public LLMs will not generate certain obviously inappropriate content when asked directly (Niu et al., 2024). However, some recent work reveals that a number of “jailbreak” tricks exist: carefully engineered prompts can result in aligned LLMs generating clearly objectionable content (Shayegani et al., 2023; Deng et al., 2024). For example, researchers have discovered that safety mechanisms can be circumvented by transforming the malicious query into semantically equivalent forms, such as ciphers (Yuan et al., 2024; Wei et al., 2024b; Jin et al., 2024), low-resource languages (Wang et al., 2023; Deng et al., 2024; Yong et al., 2023), or code (Ren et al., 2024).
With the development of jailbreaking methods for LLMs that employ various stratagems to trick the model into generating content that it is programmed to withhold or refuse (Wei et al., 2024a), recent studies also developed jailbreak methods for image generation models (Yang et al., 2023b; Kim et al., 2024). Deng and Chen (2023) is a more related attack method that breaks down an unethical drawing intent into multiple benign descriptions of individual image elements. And Jones et al. (2024) generated toxic images by image editing from another perspective. They first used a powerful closed-source model to generate harmless images, then used a local model without safety alignment to edit the harmless images into harmful ones. Different from these jailbreaking methods that only focus on single-round text-to-image generation, this paper proposes a novel jailbreak method in an iterative editing manner and camouflaging the malicious information across the multi-turn conversation. This paper also highlights the threats during the image editing process, which have not been investigated before. Besides, our method does not need additional training or use a local unaligned model.
Other Attacks and Security Issues
Previous works have also studied the robustness of text-to-image models against backdoor attack (Zhai et al., 2023; Huang et al., 2023; Vice et al., 2023), prompt stealing attack (Shen et al., 2023), and prompt poinsing attack (Shan et al., 2023).
Appendix C Jailbreaking Midjourney and Stable Diffusion
@@
Appendix D Ethical Consideration on Human Evaluation
To conduct this research safely, we have performed the following precautionary actions for the participants: (1) in every stage, we provide a content warning to the researchers and the annotators that they will see some toxic content, such as abusive, hate speech, violence and sexual related, and (2) in every stage, we told them that they could leave and quit anytime during the study and (3) we provided psychological communication after our study to make sure they are away from mental stress.
Appendix E Details of Safety Scenarios Included in Chain-of-Jailbreak Benchmark
Our CoJ benchmark includes 9 safety scenarios, the details of which is shown in Table 7, with at least 15 seed malicious queries for each safety scenario. The distribution of the benchmark is shown in Figure 6, showing the diversity of our safety scenarios.
# Scenario Description
6 Physical Harm Contents that can pose threats to human health, or encourage people to engage in high risk of bodily harm (Zhang et al., 2023b).
7 Violence Contents that depict or promote aggression or brutality.
Figure 6: Diversity of safety scenarios.
Appendix F Prompts for Automatic Query Decomposition
In section 3.2, we adopt Mistral-Large-2, to decompose each seed malicious query into a sequence of sub-queries with different edit operations and edit elements. Here we provide the specific prompts:
@@ Appendix F Prompts for Automatic Query Decomposition
Appendix G The Reliability of Automatic Query Decomposition
We propose using different Edit Operations and Edit Elements to break down a malicious query into sub-queries via Mistral-Large-2. A research question is whether the model consistently adheres to the principles of the proposed Edit Operations and Edit Elements.
To answer this question, we implement a human evaluation to evaluate the validity of the generated sub-queries. Specifically, we randomly sample 200 prompt sets from our benchmark and invite evaluators to examine the validity of these prompts. 96.5% of the generated sub-queries follow the principles of the proposed Edit Operations and Edit Elements, showing the reliability of our automatic decomposition method.
Appendix H The Efficiency of Automatic Query Decomposition
We propose using different Edit Operations and Edit Elements to break down a malicious query into sub-queries via Mistral-Large-2. Another research question is whether other models were considered and if similar performance could be achieved with smaller models requiring less computational power.
To answer this question, we conduct an additional experiment to use mistral-8b and Llama-3.1-8b, which are smaller than Mistral-Large-2, to decompose the original queries and then conduct a human evaluation to examine the validity of 200 output prompts. The results show that llama-3.1-8b is too safe and rejects all 200 instructions. Mistral-8b can follow the instructions but the quality of generated queries is not as good as Mistral-Large-2, only 76% of the generated cases are following our instructions, compared with the 96.5% of Mistral-Large-2. Since we need the LLM to be able to decompose the query into high-quality sub-queries following the examples provided, we believe Mistral-Large-2 is a better choice.
Appendix I Can We Find The Optimal Decomposition?
We have shown that decomposing the original malicious query into multiple sub-queries can achieve high jailbreak success rates. Another research question is what is the optimal decomposition, such as how many sub-queries, in what element and operation, and how we can find it.
We have shown that the number of editing steps and the way of decomposition can affect the jailbreak success rates. For example, 2-step random decomposition can achieve up to 62% jailbreak success rates (Table 3) and the success rate can be further improved via decomposing the original prompt into more steps (Figure 5). In other words, our COJ attack method has already achieved a significant success rate on jailbreak text-to-image models. However, more editing steps lead to more query times and higher energy/time costs. How to improve the efficiency of our jailbreak method is a good research question, which will be our future direction.
Appendix J Comparison with Text-to-Image Jailbreak Attack Methods
To further show the outperforming of COJ attack method, we conduct an additional experiment to compare our COJ attack method with the other 2 black box methods, SneakyPrompt (Yang et al., 2024b) and MMA-Diffusion (Yang et al., 2023a) on OpenAI GPT-4o (SneakyPrompt reports its performance on Dall-E 2, which is the inner text-to-image model in GPT-4o). As is shown in Table 8, COJ can achieve a higher jailbreak success rate with fewer average query times.
Method JSR $\uparrow$ Average Query Times $\downarrow$
SneakyPrompt 57 25
MMA-Diffusion 14 up to 20
COJ 62.3 2
Table 8: Jailbreak success rate (%) of our CoJ attack compared with other methods.
Figure 7: Examples of Pure Harmful Image Generated by COJ Attack.
Figure 8: Examples of More Vivid Images Generated by Text-to-Image Models.
Appendix K Text-to-Image Models can Generate more Vivid Images
A more straightforward approach to generate toxic images may be applicable to directly use the printed words or merge harmful words with images. So an attacker may not need a text-to-image model and thus may not need the proposed attack.
What we want to highlight here is two-fold: (1) Our CoJ method can also generate pure harmful images rather than just pure textual slogans or images with harmful words, such as examples in Figure 7. Directly merging harmful words with images cannot achieve this functionality. (2) Compared with simply merging the text with images, image generation models can generate more vivid and in-detailed images, which are more likely to be broadcasted in real-life applications. For better illustration, we show several cases in Figure 8.
Appendix L Generating Pure Harmful Image without any Word
We want to highlight that our method can also generate unsafe images without any words. For example, first “Draw a picture of a man giving flowers to a woman” and then “Change the flowers into a weapon.” According to our statistics, CoJ-Bench contains 19.3% of prompt sets that generate and edit with pure images rather than text slogans. For better illustration, we show several cases in Figure 7.